Accelerator Status Report

Дата канвертавання19.04.2016
Памер9.12 Kb.
Accelerator Status Report

Run Schedule: Aug 21, 2009 to Dec 23, 2009

Sept. 23, 2009

  • On Sept. 22, the PSS current monitor transducers of the 100/500keV spectrometer were disabled. It was believed that the current transducers had been causing beam noise. Their power supply was removed to disable the electronics. Beam current was more stable afterwards. The 100/500 keV is one of the PSS critical devices. Due to this change, the PSS injector segmentation is no longer valid, which means that the injector cannot be in beam permit with the 100/500 keV spectrometer as a PSS beam destination point.

  • On Sept. 15, beam got stuck in the injector 0L06 region again and beam transported after some valves were cycled manually. VFV0L06 was suspected to cause the problem. The valve was jumpered so that it would stay open at all times and not block beam any more.

  • Capture auto start script updated.

  • During Saturday swing shift (Sept. 12) after some injector beam line vacuum valve closed and RF 0L03 tripped, beam would not transport beyond BPM0L06. After confirming all the injector parameters were where they were supposed to be Ops suspected that a vacuum valve was not fully open even the control screen indicated all the valves were open. After cycling the valves many times beam transported.

  • Laser spot was moved yesterday, Sept. 15.

  • Wien angle is 17 degrees.


  • We continued to deliver beam with some issues: high Hall A Compton counts and ep ion chamber trips.

  • Continued commissioning of RF 1L04. Four cavities of the zone have been back online and they are stable. The remaining cavities will be back online soon.

  • Iocsb4mag went unresolved several times. Do not know the reason yet.

  • Instantaneous beam loss was not stable. It may be due to the instability of the injector.

  • When can we resume beam delivery during a vacuum recovery?

  • Pay attention after changes: LEM, ioc crash, restoration of magnets and etc.

  • Krest

  • DOG2E and DOG3E tripped off several times and EES fixed them.

  • RF separators had cooling problems and kept tripping. They were acid flushed yesterday.

  • We have tightened up ion chamber trip levels in Hall A and Hall C. We may have to relax the trip levels if we have more trips but make sure beam transports well when adjusting them.

Hall A---E05-109; pass 3; energy 3.49 GeV

  • Hall A target or ep ion chamber would trip off beam when Hall A dithered its beam. It was found that the root cause was the FFB. The transition of the FFB on/off would cause the trips. Iocse9 was power cycled, the FFB sever re-started, the FFB re-calibrated, the BMP gain switched to manual mode for both the FFB system of Halls A and C. after the above actions the ion chamber trips stopped.

  • Hall A Compton counts were too high. Several hours were spent to troubleshoot the problem. Basically we cleaned up the entire machine from the injector to Hall A without good results. It turned out that the Compton test signal was left on. The Compton counts were good after the test signal was turned off.

  • BPM1C12 saturation.

  • With optics on-call support many hours were spent to lower and stabilize Hall A Compton counts. More time is needed to troubleshoot the problem.

  • Hall A Compton ion chamber sometimes trips off even with on beam in the hall. We may have to raise the trip level a bit. We are having problems raising the Compton trip level. It would not go higher than 1000 and Lahti is looking into it.

Hall B ---Eg1-DVCS; pass 5; energy 5.764 GeV

  • Hall B is happy with its beam.

  • Hall B will be down for ten days to change to a new experiment.

  • Monday Hall B reported that we sent too much beam to their during the beam checkout after a down. Hall B has a polarized target. The polarization of its target could be spoiled by current higher than 10nA. If we need to send tune beam to the hall line we can let the hall move out its target or send beam to its Tagger dump.

  • Hall B orbit lock was fixed and has been working.

Hall C---HKS; pass 2; energy 2,344 GeV

  • During Sept. 17 swing shift FFB slow lock steered beam away from the target positions a few times. S. Higgins investigated the problem but did not find anything yet.

  • Hall C fast raster does not work. Be careful when the hall changes its target.

  • Hall C SLI indicates the energy spread higher than required (5E-5).

  • Hall C beam sizes are larger than required (50um x 100um).

База данных защищена авторским правом © 2016
звярнуцца да адміністрацыі

    Галоўная старонка